Two angles in a triangle are equal and their sum is equal to the third angle in the triangle. What are the measures of each of the three interior angles?

Found 2 solutions by sudhanshu_kmr, Alan3354:
Answer by sudhanshu_kmr(1152)


Assume angles are x then third is x+x = 2x.
sum of angles in a triangle = 180
=> x + x + 2x = 180
=> 4x = 180
=> x = 180/4 = 45

so, angles are 45,45 and 90.

Answer by Alan3354(69443)

45, 45, 90 degs